# Knowledge Mechanisms in Large Language Models: A Survey and Perspective

### v2

@@ 3.4 Comparison of Different Mechanism Analysis Methods
Each method is suitable for different scenarios. Specifically, the Modular Region Hypothesis can be analyzed using either Observation-based or Intervention-based methods. In contrast, the Connection Hypothesis, which examines inter-regional connectivity, generally necessitates Intervention-based methods. However, the results of knowledge mechanism analysis depend heavily on different methods and are sensitive to evaluation metrics and implementation details. Hence, Huang et al. (2024b) propose a dataset, RAVEL, to quantify the comparisons between a variety of existing interpretability methods. They suggest that methods with supervision are better than methods with unsupervised featurizers. Later, Zhang and Nanda (2023) further systematically examine the impact of methodological details in intervention-based methods. For corrupted run, they recommend Symmetric Token Replacement (e.g., “The Eiffel Tower”$\rightarrow$“The Colosseum”) Sharma et al. (2024); Wang et al. (2023c); Vig et al. (2020) instead of Gaussian Noising Meng et al. (2022), which disrupts the model’s internal mechanisms. Besides, they advocate using the logit lens over probes for evaluation metric $E$ due to its fine-grained control over localization outcomes.
@@ 4.1 Individual Evolution
In the pre-training phase, LLMs start as blank slates, facilitating easier acquisition for new knowledge. Consequently, numerous experiments demonstrate that LLMs accumulate vast amounts of knowledge during this stage Cao et al. (2024b); Zhou et al. (2023a); Kaddour et al. (2023); Naveed et al. (2023); Singhal et al. (2022). Akyürek et al. (2022) delve further into identifying which training examples are instrumental in endowing LLMs with specific knowledge. However, contradictions during the pre-training stage may induce conflicts among internal parametric knowledge. On the one hand, the false and contradictory information in training corpus propagate and contaminate related memories in LLMs via semantic diffusion, introducing broader detrimental effects beyond direct impacts Bian et al. (2023). On the other hand, LLMs tend to prioritize memorizing more frequent and challenging facts, which can result in subsequent facts overwriting prior memorization, significantly hindering the memorization of low-frequency facts Lu et al. (2024). In other words, LLMs struggle with balancing and integrating both low and high-frequency knowledge.
@@
After pre-training, LLMs are anticipated to refresh their internal knowledge to keep pace with the evolving world during post-training stage. Although LLMs seem to absorb new knowledge through continued learning, follow user instructions via instruct tuning Zhang et al. (2023c), and align with human values through alignment tuning Ziegler et al. (2019), Ji et al. (2024a) have noted that LLMs intrinsically resist alignment during the post-training phase. In other words, LLMs tend to learn factual knowledge through pre-training, whereas fine-tuning  [^4] Fine-tuning includes instruct tuning and alignment tuning Zhao et al. (2023b). teaches them to utilize it more efficiently Gekhman et al. (2024); Zhou et al. (2023a); Ovadia et al. (2024). Ren et al. (2024) also posit that instruction tuning is a form of self-alignment with existing internal knowledge rather than a process of learning new information. We conjecture that the debate on whether these processes truly introduce new knowledge stems from information conflicts. For example, the conflict between outdated information within LLMs and new external knowledge exacerbates their difficulty in learning new information. To mitigate information conflicts, Ni et al. (2023) propose first forgetting old knowledge then learning new knowledge. Another technique, retrieval-augmented generation (RAG) Huang and Huang (2024), while avoiding conflicts within internal parameters, still needs to manage conflicts between retrieved external information and LLMs’ internal knowledge Xu et al. (2024b). RAG also attempt to efficiently and effectively integrate new knowledge across passages or documents using multiple retrieval Yang et al. (2024a) and hippocampal indexing Gutiérrez et al. (2024). Besides, editing technologies, including knowledge and representation editing, exhibit promising potential for knowledge addition, modification, and erasure. Specifically, knowledge editing Meng et al. (2022); Mitchell et al. (2022); Cao et al. (2021b); Zhang et al. (2024a); Wang et al. (2023d); Mazzia et al. (2023) aims to selectively modify model parameters responsible for specific knowledge retention, while representation editing Zou et al. (2023); Wu et al. (2024) adjusts the model’s conceptualization of knowledge to revise the stored knowledge within LLMs. Note that the other strategy for knowledge editing adds external parameters or memory banks for new knowledge while preserving models’ parameters. We also provide the comparison of the above methods in §A for better understanding.
@@ 4.2 Group Evolution
Besides individual learning, social interaction plays a pivotal role in the acquisition of new knowledge and is a key driver of human societal development Baucal et al. (2014). LLMs, also known as agents, collaborate to accomplish complex tasks during group evolution, each bearing unique knowledge that may sometimes contradict each other. Therefore, contrary to individual evolution, group evolution encounters intensified conflicts, such as conflicts in specialized expertise among agents, competing interests, cultural disparities, moral dilemmas, and others. To achieve consensus and resolve conflicts, agents must first clarify their own and others’ goals (beliefs) through internal representations in models Zhu et al. (2024); Zou et al. (2023). Agents then discuss, debate, and reflect on shared knowledge through various communication methods Soltoggio et al. (2024), e.g., prompt instructions, task and agent descriptions, parameter signals (activation and gradient), and representations of models. However, conformity of agents, which tends to believe the majority’s incorrect answers rather than maintaining their own, hinders conflict resolution during group evolution Zhang et al. (2023a). Note that the group also struggles with automating moral decision-making when facing moral conflicts. Specifically, agents in the group miss ground truth for moral “correctness” and encounter dilemmas due to changes in moral norms over time Hagendorff and Danks (2023). Generally, when, what, and how to share knowledge in the communication process to maximize learning efficiency and long-term expectations are still open questions in group evolution.
@@ 5.2 Trustworthy LLMs
Numerous studies investigate the underlying causes of security risks. In particular, Wei et al. (2023) delve into the safety of LLM and reveal that the success of jailbreak is mainly due to the distribution discrepancies between malicious attacks and training data. Geva et al. (2022b) and Wang et al. (2024b) further discover that some parameters within LLMs, called toxic regions, are intrinsically tied to the generation of toxic content. Ji et al. (2024a) even conjecture that LLMs resist alignment. Therefore, traditional aligned methods, DPO Rafailov et al. (2023) and SFT, seem to merely bypass toxic regions Lee et al. (2024); Wang et al. (2024b), making them susceptible to other jailbreak attacks Zhang et al. (2023d).
@@
Inspired by the knowledge mechanism analysis in LLMs, a promising trustworthy strategy may be designing architecture and training process during the pre-training phase to encouraging modularity Liu et al. (2024c, 2023b), sparsity Chughtai et al. (2023), and monosemanticity Bricken et al. (2023); Jermyn et al. (2022), which make the reverse engineering process more tractable Jermyn et al. (2022); Bricken et al. (2023); Liu et al. (2024c); Tamkin et al. (2023). Yet, maintaining sparsity for a vast amount of world knowledge requires substantial resources, and whether monosemantic architecture can support advanced intelligence remains elusive. Besides, machine unlearning Nguyen et al. (2022); Tian et al. (2024) aims to forget privacy or toxic information learned by LLMs. However, these unlearning methods suffer overfitting, forgetting something valuable due to the difficulty of disentangling verbatim memorization and general capabilities Huang et al. (2024c). Another alternative technique is knowledge editing, precisely modifying LLMs using few instances during the post-training stage Mazzia et al. (2023); Yao et al. (2023); Wang et al. (2023d); Hase et al. (2024); Qian et al. (2024b). Extensive experiments demonstrate that knowledge editing has the potential to detoxify LLMs Yan et al. (2024). Specifically, Wu et al. (2023) and Geva et al. (2022b) deactivate the neurons related to privacy information and toxic tokens, respectively. Wang et al. (2024b) identify and then erases toxic regions in LLMs. However, knowledge editing also introduces side effects, such as the inability of the modified knowledge to generalize to multi-hop tasks Zhong et al. (2023); Li et al. (2023d); Cohen et al. (2023); Kong et al. (2024) and the potential to impair the model’s general capabilities Gu et al. (2024); Qin et al. (2024). Therefore, recent efforts focus on representation editing instead of editing parameters in knowledge editing Zou et al. (2023); Turner et al. (2023); Zhou et al. (2023b); Zhu et al. (2024). These representations (hidden states) within LLMs can track and address a wide range of safety-relevant problems, including honesty, harmlessness, and power seeking. Later, Wu et al. (2024) develop a family of representation finetuning methods to update new knowledge. Zou et al. (2024) propose circuit-breaking Li et al. (2023b), directly controlling the representations that are responsible for harmful outputs. However, these representation editing strategies require meticulous hyperparameter tuning for each task. More efficient optimization methods are needed to align with computational or temporal constraints.
@@ 6.1 What Knowledge Have LLMs Learned?
Critics question whether LLMs truly have knowledge or if they are merely mimicking Schwarzschild et al. (2024), akin to the “Stochastic Parro” Bender et al. (2021) and “Clever Hans” Shapira et al. (2024). We first review the doubts from the following three levels through observation phenomena: 1) Memorization: LLMs primarily rely on positional information over semantic understanding Li et al. (2022) to predict answers. Additionally, LLMs may generate different answers for the same question due to different expressions. 2) Comprehension and application: Allen-Zhu and Li (2023b) argue that LLMs hardly efficiently apply knowledge from pre-training data, even when such knowledge is perfectly stored and fully extracted from LLMs. Therefore, LLMs struggle with diverse reasoning tasks Nezhurina et al. (2024); Gutiérrez et al. (2024) as well as the reversal curse Berglund et al. (2023). Besides, LLMs are not yet able to reliably act as text world simulators and encounter difficulties with planning Wang et al. (2024d). 3) Creation: Although LLMs are capable of generating new terms, their quality often falls below that created by humans Raiola (2023). Even though LLMs possess knowledge, some critics argue that current analysis methods may only explain low-level co-occurrence patterns, not internal mechanisms. The primary criticism asserts that the components responsible for certain types of knowledge in LLM fail to perform effectively in practical applications Hase et al. (2023). In addition, the components responsible for specific knowledge within LLMs vary under different methods. For these criticisms, Chen et al. (2024f, d) propose degenerate neurons and posit that different degenerate components indeed independently express a fact. Chen et al. (2024e) delineate the differences in the mechanisms of knowledge storage and representation, proposing the Query Localization Assumption to response these controversies. Zhu and Li (2023) further observe that knowledge may be memorized but not extracted due to the knowledge not being sufficiently augmented (e.g., through paraphrasing, sentence shuffling) during pretraining. Hence, rewriting the training data to provide knowledge augmentation and incorporating more instruction fine-tuning data in the pre-training stage can effectively alleviate the above challenges and criticisms.
@@
Despite considerable criticism, the mainstream view is that current LLMs may possess basic world knowledge but hardly master principle knowledge of underlying rules for reasoning and creativity. In other words, LLMs master knowledge memorization (discussed in §3.1). Although LLMs possess the foundational ability to comprehend and apply knowledge (discussed in §3.2), they still struggle with reasoning and planning in complex tasks due to the fragility of knowledge in LLMs (elaborated in §6.2). These reasoning and planning abilities usually require to be induced through techniques such as ICL and CoT. Unfortunately, current LLMs are nearly incapable of creation due to the architectural limitations (discussed in §3.3). Therefore, some scholars explore various architectural choices (e.g., Mamba Gu and Dao (2023)) and training procedures. Besides, recent research attempts to manipulate neurons, knowledge circuits, or representations Allen-Zhu and Li (2023b); Zou et al. (2023); Wu et al. (2024); Li et al. (2023a) to explore more knowledge and awaken the reasoning and planning capabilities of LLMs.
@@ 6.2 Why Is Learned Knowledge Fragile?
The knowledge learned by LLMs is fragile, leading to challenges in application including hallucination, knowledge conflicts, and secure risk  [^8] The secure risk is elaborated in §5.2.. Hallucination denotes content generated by LLMs that diverges from real-world facts or inputs Huang et al. (2023b); Xu et al. (2024d); Farquhar et al. (2024); Chen et al. (2024c). On the one hand, factuality hallucination underscores the disparity between generated content and real-world knowledge. On the other hand, faithfulness hallucination describes the departure of generated content from user instructions or input context, as well as the coherence maintained within the generated content. Knowledge Conflict inherently denotes inconsistencies in knowledge Xu et al. (2024b); Kortukov et al. (2024). On the one hand, internal memory conflicts within the model cause LLMs to exhibit unpredictable behaviors and generate differing results to inputs which are semantically equivalent but syntactically distinct Xu et al. (2024b); Wang et al. (2023a); Feng et al. (2023b); Raj et al. (2022). On the other hand, context-memory conflict emerges when external context knowledge contradicts internal parametric knowledge Xu et al. (2024b); Mallen et al. (2023).
@@
We posit that these challenges mainly derive from improper learning data. Specifically, hallucination is introduced by data Kang and Choi (2023); Weng (2024), heightened during the pre-training Brown et al. (2020); Chiang and Cholak (2022), alignment Azaria and Mitchell (2023); Ouyang et al. (2022), and deficiencies in decoding strategies Fan et al. (2018); Chuang et al. (2023); Shi et al. (2023). Internal memory conflict can be attributed to training corpus bias Wang et al. (2023b), and exacerbated by decoding strategies Lee et al. (2022b) and knowledge editing. Context-memory conflict arises mainly from the absence of accurate knowledge during training, necessitating retrieval from databases and the Web. Wang et al. (2024a) and Antoniades et al. (2024) further delve into the mechanism between parametric knowledge and learning data, demonstrate that training data distribution qualitatively influences generalization behavior Jiang et al. (2024a) and how LLMs encapsulate or enhance knowledge related to reasoning into specific components, e.g., bridge layer Wang et al. (2024a). Similar mechanism analysis also supports the above conclusion, indicating that hallucinations arise from a lack of mover heads Yao et al. (2024); Yu et al. (2024b), while knowledge conflicts stem from circuit competition failure in the last few layers Lv et al. (2024); Merullo et al. (2023b); Hase et al. (2023); Ju et al. (2024); Jin et al. (2024b). Additionally, data quantity is crucial for knowledge robustness. Specifically, LLMs can systematically learn comprehensive understandings of the world from extensive datasets, while little data during post-training stage may compromise the robustness of knowledge representation. This assumption is confirmed by numerous failures of post-training. For example, SFT exacerbates hallucinations Gekhman et al. (2024); Kang et al. (2024), and knowledge editing amplifies knowledge conflicts Li et al. (2023d); Yang et al. (2024c). Note that safety issues usually caused by the distribution of unseen data (adversarial input) Wei et al. (2023); Li et al. (2024b), which is elaborated in §5.2.
@@
Remarks: The knowledge fragility of LLMs may be attributed to misuse Ji et al. (2024b), but improper learning might be the fundamental and primary cause.

### v3

@@ 3.4 Comparison of Different Mechanism Analysis Methods
Each method is suitable for different scenarios. Specifically, the Modular Region Hypothesis can be analyzed using either Observation-based or Intervention-based methods. In contrast, the Connection Hypothesis, which examines inter-regional connectivity, generally necessitates Intervention-based methods. However, the results of knowledge mechanism analysis depend heavily on different methods and are sensitive to evaluation metrics and implementation details Schwettmann et al. (2023b). Hence, Huang et al. (2024b) propose a dataset, RAVEL, to quantify the comparisons between a variety of existing interpretability methods. They suggest that methods with supervision are better than methods with unsupervised featurizers. Later, Zhang and Nanda (2023) further systematically examine the impact of methodological details in intervention-based methods. For corrupted run, they recommend Symmetric Token Replacement (e.g., “The Eiffel Tower”$\rightarrow$“The Colosseum”) Sharma et al. (2024); Vig et al. (2020) instead of Gaussian Noising Meng et al. (2022), which disrupts the model’s internal mechanisms. For metric $E$, both logit lens and probe can be employed to trace factual knowledge Meng et al. (2022), where the target output is typically few tokens. In this scenario, Zhang and Nanda (2023) advocate using the logit lens over probes for evaluation metric $E$ due to its fine-grained control over localization outcomes. Moreover, probe is capable of exploring abstract knowledge and abilities, such as theory of mind or mental states Zhu et al. (2024); Ye et al. (2024); Jin (2024), where the target output requires multiple tokens to express. Jin (2024) suggest that deeper probes are more (generally) more accurate.
@@ 4.1 Individual Evolution
In the pre-training phase, LLMs start as blank slates, facilitating easier acquisition for new knowledge Allen-Zhu and Li (2024). Consequently, numerous experiments demonstrate that LLMs accumulate vast amounts of knowledge during this stage Cao et al. (2024b); Zhou et al. (2023a); Kaddour et al. (2023); Naveed et al. (2023); Singhal et al. (2022). Akyürek et al. (2022) delve further into identifying which training examples are instrumental in endowing LLMs with specific knowledge. However, contradictions during the pre-training stage may induce conflicts among internal parametric knowledge. On the one hand, the false and contradictory information in training corpus propagate and contaminate related memories in LLMs via semantic diffusion, introducing broader detrimental effects beyond direct impacts Bian et al. (2023). On the other hand, LLMs tend to prioritize memorizing more frequent and challenging facts, which can result in subsequent facts overwriting prior memorization, significantly hindering the memorization of low-frequency facts Lu et al. (2024). In other words, LLMs struggle with balancing and integrating both low and high-frequency knowledge.
@@
After pre-training, LLMs are anticipated to refresh their internal knowledge to keep pace with the evolving world during post-training stage. Although LLMs seem to absorb new knowledge through continued learning, follow user instructions via instruct tuning Zhang et al. (2023c), and align with human values through alignment tuning Ziegler et al. (2019), Ji et al. (2024) have noted that LLMs intrinsically resist alignment during the post-training phase. In other words, LLMs tend to learn factual knowledge through pre-training, whereas fine-tuning  [^4] Fine-tuning includes instruct tuning and alignment tuning Zhao et al. (2023b). teaches them to utilize it more efficiently Gekhman et al. (2024); Zhou et al. (2023a); Ovadia et al. (2024). Ren et al. (2024a) also posit that instruction tuning is a form of self-alignment with existing internal knowledge rather than a process of learning new information. We conjecture that the debate on whether these processes truly introduce new knowledge stems from information conflicts. For example, the conflict between outdated information within LLMs and new external knowledge exacerbates their difficulty in learning new information. To mitigate information conflicts, Ni et al. (2023) propose first forgetting old knowledge then learning new knowledge. Another technique, retrieval-augmented generation (RAG) Huang and Huang (2024), while avoiding conflicts within internal parameters, still needs to manage conflicts between retrieved external information and LLMs’ internal knowledge Xu et al. (2024b). RAG also attempt to efficiently and effectively integrate new knowledge across passages or documents using multiple retrieval Yang et al. (2024a) and hippocampal indexing Gutiérrez et al. (2024). Besides, editing technologies, including knowledge and representation editing, exhibit promising potential for knowledge addition, modification, and erasure. Specifically, knowledge editing Meng et al. (2022); Mitchell et al. (2022); Cao et al. (2021b); Zhang et al. (2024a); Wang et al. (2023d); Mazzia et al. (2023) aims to selectively modify model parameters responsible for specific knowledge retention, while representation editing Zou et al. (2023); Wu et al. (2024) adjusts the model’s conceptualization of knowledge to revise the stored knowledge within LLMs. Note that the other strategy for knowledge editing adds external parameters or memory banks for new knowledge while preserving models’ parameters. We also provide the comparison of the above methods in §A for better understanding.
@@ 4.2 Group Evolution
Besides individual learning, social interaction plays a pivotal role in the acquisition of new knowledge and is a key driver of human societal development Baucal et al. (2014); Levine et al. (1993). LLMs, also known as agents, collaborate to accomplish complex tasks during group evolution, each bearing unique knowledge that may sometimes contradict each other. Therefore, contrary to individual evolution, group evolution encounters intensified conflicts, such as conflicts in specialized expertise among agents, competing interests, cultural disparities, moral dilemmas, and others. To achieve consensus and resolve conflicts, agents must first clarify their own and others’ goals (beliefs) through internal representations in models Zhu et al. (2024); Zou et al. (2023). Agents then discuss, debate, and reflect on shared knowledge through various communication methods Chan et al. (2024); Smit et al. (2024); Li et al. (2024e); Soltoggio et al. (2024), e.g., prompt instructions, task and agent descriptions, parameter signals (activation and gradient), and representations of models. However, conformity of agents, which tends to believe the majority’s incorrect answers rather than maintaining their own, hinders conflict resolution during group evolution Zhang et al. (2023a); Ma et al. (2024). Note that the group also struggles with automating moral decision-making when facing moral conflicts. Specifically, agents in the group miss ground truth for moral “correctness” and encounter dilemmas due to changes in moral norms over time Hagendorff and Danks (2023). Generally, when, what, and how to share knowledge in the communication process to maximize learning efficiency and long-term expectations are still open questions in group evolution.
@@ 5.2 Trustworthy LLMs
Numerous studies investigate the underlying causes of security risks Reuel et al. (2024); Ren et al. (2024b); Li et al. (2024a); Bengio (2024); Bengio et al. (2024); Dalrymple et al. (2024). In particular, Wei et al. (2023) delve into the safety of LLM and reveal that the success of jailbreak is mainly due to the distribution discrepancies between malicious attacks and training data. Geva et al. (2022b) and Wang et al. (2024b) further discover that some parameters within LLMs, called toxic regions, are intrinsically tied to the generation of toxic content. Ji et al. (2024) even conjecture that LLMs resist alignment. Therefore, traditional aligned methods, DPO Rafailov et al. (2023) and SFT, seem to merely bypass toxic regions Lee et al. (2024); Wang et al. (2024b), making them susceptible to other jailbreak attacks Zhang et al. (2023d).
@@
Inspired by the knowledge mechanism analysis in LLMs, a promising trustworthy strategy may be designing architecture and training process during the pre-training phase to encouraging modularity Liu et al. (2024c, 2023b), sparsity Chughtai et al. (2023), and monosemanticity Bricken et al. (2023); Jermyn et al. (2022), which make the reverse engineering process more tractable Jermyn et al. (2022); Bricken et al. (2023); Liu et al. (2024c); Tamkin et al. (2023). Yet, maintaining sparsity for a vast amount of world knowledge requires substantial resources, and whether monosemantic architecture can support advanced intelligence remains elusive. Besides, machine unlearning Nguyen et al. (2022); Tian et al. (2024); Yao et al. (2023a) aims to forget privacy or toxic information learned by LLMs. However, these unlearning methods suffer overfitting, forgetting something valuable due to the difficulty of disentangling verbatim memorization and general capabilities Huang et al. (2024c); Blanco-Justicia et al. (2024). Another alternative technique is knowledge editing, precisely modifying LLMs using few instances during the post-training stage Mazzia et al. (2023); Yao et al. (2023b); Wang et al. (2023d); Hase et al. (2024); Qian et al. (2024b). Extensive experiments demonstrate that knowledge editing has the potential to detoxify LLMs Yan et al. (2024). Specifically, Wu et al. (2023a) and Geva et al. (2022b) deactivate the neurons related to privacy information and toxic tokens, respectively. Wang et al. (2024b) identify and then erases toxic regions in LLMs. However, knowledge editing also introduces side effects, such as the inability of the modified knowledge to generalize to multi-hop tasks Zhong et al. (2023); Li et al. (2023d); Cohen et al. (2023); Kong et al. (2024) and the potential to impair the model’s general capabilities Gu et al. (2024); Qin et al. (2024). Therefore, recent efforts focus on representation editing instead of editing parameters in knowledge editing Zou et al. (2023); Turner et al. (2023); Zhou et al. (2023b); Zhu et al. (2024). These representations (hidden states) within LLMs can trace and address a wide range of safety-relevant problems, including honesty, harmlessness, and power seeking. Later, Wu et al. (2024) develop a family of representation finetuning methods to update new knowledge. Zou et al. (2024) propose circuit-breaking Li et al. (2023b), directly controlling the representations that are responsible for harmful outputs. However, these representation editing strategies require meticulous hyperparameter tuning for each task. More efficient optimization methods are needed to align with computational or temporal constraints.
@@ 6.1 What Knowledge Have LLMs Learned?
Critics question whether LLMs truly have knowledge or if they are merely mimicking Schwarzschild et al. (2024), akin to the “Stochastic Parro” Bender et al. (2021) and “Clever Hans” Shapira et al. (2024). We first review the doubts from the following three levels through observation phenomena: 1) Memorization: LLMs primarily rely on positional information over semantic understanding Li et al. (2022) to predict answers. Additionally, LLMs may generate different answers for the same question due to different expressions. 2) Comprehension and application: Allen-Zhu and Li (2023b) argue that LLMs hardly efficiently apply knowledge from pre-training data, even when such knowledge is perfectly stored and fully extracted from LLMs. Therefore, LLMs struggle with various reasoning tasks Wu et al. (2023b); Nezhurina et al. (2024); Gutiérrez et al. (2024) as well as the reversal curse Berglund et al. (2023). Besides, LLMs are not yet able to reliably act as text world simulators and encounter difficulties with planning Wang et al. (2024d). 3) Creation: Although LLMs are capable of generating new terms, their quality often falls below that created by humans Raiola (2023). Even though LLMs possess knowledge, some critics argue that current analysis methods may only explain low-level co-occurrence patterns, not internal mechanisms. The primary criticism asserts that the components responsible for certain types of knowledge in LLM fail to perform effectively in practical applications Hase et al. (2023). In addition, the components responsible for specific knowledge within LLMs vary under different methods. For these criticisms, Chen et al. (2024f, d) propose degenerate neurons and posit that different degenerate components indeed independently express a fact. Chen et al. (2024e) delineate the differences in the mechanisms of knowledge storage and representation, proposing the Query Localization Assumption to response these controversies. Zhu and Li (2023) further observe that knowledge may be memorized but not extracted due to the knowledge not being sufficiently augmented (e.g., through paraphrasing, sentence shuffling) during pretraining. Hence, rewriting the training data to provide knowledge augmentation and incorporating more instruction fine-tuning data in the pre-training stage can effectively alleviate the above challenges and criticisms.
@@
Despite considerable criticism, the mainstream view Didolkar et al. (2024); Jin and Rinard ; Jin (2024) is that current LLMs may possess basic world knowledge via memorization but hardly master underlying principles for reasoning and creativity. In other words, LLMs master basic knowledge via memorization (discussed in §3.1). Although LLMs possess the foundational ability to comprehend and apply knowledge (discussed in §3.2), exhibiting plausible and impressive reasoning capabilities. Current LLMs still struggle with reasoning and planning in complex tasks due to the fragility of knowledge in LLMs (elaborated in §6.2). These reasoning and planning abilities usually require to be induced through techniques such as ICL and CoT. Unfortunately, current LLMs are nearly incapable of creation due to the architectural limitations (discussed in §3.3). Therefore, some scholars explore various architectural choices (e.g., Mamba Gu and Dao (2023)) and training procedures. Besides, recent research attempts to manipulate neurons, knowledge circuits, or representations Allen-Zhu and Li (2023b); Zou et al. (2023); Wu et al. (2024); Li et al. (2023a) to explore more knowledge and awaken the reasoning and planning capabilities of LLMs.
@@ 6.2 Why Is Learned Knowledge Fragile?
The knowledge learned by LLMs is fragile, leading to challenges in application including hallucination, knowledge conflicts, failed reasoning, and safety risk  [^8] The secure risk is elaborated in §5.2.. Hallucination denotes content generated by LLMs that diverges from real-world facts or inputs Huang et al. (2023b); Xu et al. (2024d); Farquhar et al. (2024); Chen et al. (2024c). On the one hand, factuality hallucination underscores the disparity between generated content and real-world knowledge. On the other hand, faithfulness hallucination describes the departure of generated content from user instructions or input context, as well as the coherence maintained within the generated content. Knowledge Conflict inherently denotes inconsistencies in knowledge Xu et al. (2024b); Kortukov et al. (2024). On the one hand, internal memory conflicts within the model cause LLMs to exhibit unpredictable behaviors and generate differing results to inputs which are semantically equivalent but syntactically distinct Xu et al. (2024b); Wang et al. (2023a); Feng et al. (2023b); Raj et al. (2022). On the other hand, context-memory conflict emerges when external context knowledge contradicts internal parametric knowledge Xu et al. (2024b); Mallen et al. (2023).
@@
We posit that these challenges mainly derive from improper learning data. Specifically, hallucination is introduced by data Kang and Choi (2023); Weng (2024); Zhang et al. (2024c), heightened during the pre-training Brown et al. (2020); Chiang and Cholak (2022), alignment Azaria and Mitchell (2023); Ouyang et al. (2022), and deficiencies in decoding strategies Fan et al. (2018); Chuang et al. (2023); Shi et al. (2023). Internal memory conflict can be attributed to training corpus bias Wang et al. (2023b), and exacerbated by decoding strategies Lee et al. (2022b) and knowledge editing. Context-memory conflict arises mainly from the absence of accurate knowledge during training, necessitating retrieval from databases and the Web. Failed reasoning usually arises from improper data distribution. Specifically, knowledge may be memorized but not extractable or applicable without sufficient augmentation (e.g., through paraphrasing, sentence shuffling) during pre-training Zhu and Li (2023). Antoniades et al. (2024) also delve into the mechanism between parametric knowledge and learning data, demonstrate that training data distribution qualitatively influences generalization behavior Jiang et al. (2024a). Wang et al. (2024a) further suggest that improper data distribution in the corpus causes LLMs to lack essential reasoning components, such as the bridge layer for two-hop reasoning. Similar mechanism analysis also supports the above conclusion, indicating that hallucinations arise from a lack of mover heads Yao et al. (2024); Yu et al. (2024b), while knowledge conflicts stem from circuit competition failure in the last few layers Lv et al. (2024); Merullo et al. (2023b); Hase et al. (2023); Ju et al. (2024); Jin et al. (2024b). Additionally, data quantity is crucial for knowledge robustness. Specifically, LLMs can systematically learn comprehensive understandings of the world from extensive datasets, while little data during post-training stage may compromise the robustness of knowledge representation. This assumption is confirmed by numerous failures of post-training. For example, SFT exacerbates hallucinations Gekhman et al. (2024); Kang et al. (2024), and knowledge editing amplifies knowledge conflicts Li et al. (2023d); Yang et al. (2024c). Note that safety issues usually caused by the distribution of unseen data (adversarial input) Wei et al. (2023); Li et al. (2024c), which is elaborated in §5.2.
@@
Remarks: Improper learning caused by data distribution and quantity might be the fundamental and primary cause.